OpenStudy (anonymous):

-3 2/5 + b= 8 1/5 @iGreen

OpenStudy (igreen):

We gotta add \(3 \dfrac{2}{5}\) to both sides. \(8 \dfrac{1}{5} + 3 \dfrac{2}{5}\) These have the same denominator..so we just have to switch then both into improper fractions.

OpenStudy (igreen):

Can you switch them both into improper fractions?

OpenStudy (anonymous):

41/5 + 17/5

OpenStudy (igreen):

Yep, you got it! Now add them.

OpenStudy (igreen):

41 + 17 = ?

OpenStudy (anonymous):

58/5

OpenStudy (igreen):

Yep, you got it..now convert back to mixed number.

OpenStudy (anonymous):

11 3/5

OpenStudy (igreen):

Yep! :D Outstanding work.

OpenStudy (anonymous):

2 more?

OpenStudy (igreen):

Yep, I think I can do that..

OpenStudy (anonymous):

-7/3n=-42

OpenStudy (igreen):

You have to divide -7/3 to both sides, or multiply by its reciprocal.. \(-7/3n = -42\) \(n = -42 \times -3/7\)

OpenStudy (igreen):

Multiply that, what do you get?

OpenStudy (anonymous):

donno

OpenStudy (anonymous):

notsure

OpenStudy (igreen):

\(-\dfrac{42}{1} \times -\dfrac{3}{7} \rightarrow \dfrac{-42 \times -3}{-1 \times -7}\) Can you simplify that?

OpenStudy (anonymous):

-126/7

OpenStudy (igreen):

Yep, now divide that.

OpenStudy (igreen):

Actually it would be 126/7

OpenStudy (anonymous):

okay

OpenStudy (igreen):

Can you divide that?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

10 2/3 - 8 3/4

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen